OpenStudy (anonymous):

Solve using synthetic division: (x^3+x^2-9x-9) divided by (x+1)

myininaya (myininaya):

x+1=0 => x=-1 1x^3+1x^2-9x-9 -1 | 1 1 -9 -9 | -1 0 9 ========== 1 0 -9 | 0 \[\frac{x^3+x^2-9x-9}{x+1}=x^2+0x-9=x^2-9\]

OpenStudy (anonymous):

how did you write out such pretty formulas?

myininaya (myininaya):

oh sorry moses i didn't mean to leave you hanging

myininaya (myininaya):

do frac{ }{ } in the equation editor below

myininaya (myininaya):

frac{ something}{ something} will give \[\frac{ something}{ something} \]

myininaya (myininaya):

\[\int\limits_{a}^{b}f(x) dx\]

OpenStudy (anonymous):

{hi}{bi}

myininaya (myininaya):

see the equation button below

OpenStudy (anonymous):

frac{hi}{bi}

OpenStudy (anonymous):

\[\frac{hi}{bi} \]

OpenStudy (anonymous):

ah i see. is that latex?

myininaya (myininaya):

yay!

myininaya (myininaya):

yes

OpenStudy (anonymous):

that is so sexy!

myininaya (myininaya):

you don't have to go to equation to type there is another way

OpenStudy (anonymous):

teach me latex!!!!!

myininaya (myininaya):